OpenStudy (edwardcullen2017):

what is Y= (7*7-4)/(17-2)

hartnn (hartnn):

what is 17-2=?

hartnn (hartnn):

ok, what is 7*7=?

OpenStudy (edwardcullen2017):

15 and then 49

hartnn (hartnn):

ok, so numerator = 49-4=? and denominator = 15

OpenStudy (edwardcullen2017):

oh sorry i made a simple mistake..im dumb...thanks

hartnn (hartnn):

did u get 3 as answer ?

hartnn (hartnn):

45/15=3

OpenStudy (edwardcullen2017):

yeah i did

hartnn (hartnn):

good ^_^